A washer and a dryer cost $560 combined. The washer costs $90 less than the dryer. What is the cost of the dryer?
Answer by ikleyn(52777)

.
A washer and a dryer cost $560 combined. The washer costs $90 less than the dryer.
What is the cost of the dryer?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let X be the cost of the dryer, in dollars.

Then the cost of the washer  is (X-90) dollars.


Together they cost $560

    X + (X-90) = 560

    2X = 560 + 90

    2X = 650

     X = 650/2 = 325.


ANSWER.  The dryer costs $325.

Solved.

-------------------

To see million of other similar and different problems solved,  look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.